# Liu and Zhang

### v2

@@ 4 Conclusion
In this paper, we generalize the PCL framework for RMAB with finite state spaces to countable state spaces and propose an efficient algorithm to test the PCL-indexability and compute the Whittle index function for the partially observable RMAB with general observation models. As the number of belief states increases rapidly with time, the classic dynamic programming analysis for the RMAB problem suffers from the curse of dimensionality. Based on the achievable region and conservation law analysis studied in the literature for various RMAB problems, this paper is the first to extend the previous theoretical results in this direction to infinite-state scenarios by taking advantage of the theory of infinite-dimensional linear programming. Under a loose regularity assumption, we prove that the PCL-indexability is a sufficient condition for the Whittle indexability. Last, by introducing the concept of approximate state space, we build an algorithm to calculate the approximate Whittle index of each state using a finite-state adaptive greedy algorithm. Numerical results verified the strong performance of the Whittle index policy in our model.
Although we have quite a few theoretical analysis and confirmed the effectiveness of the Whittle index algorithm, there are still some problems to be solved in the future. An important one is that the concept of approximate state space is formed on the basis of a large number of numerical simulations, and we have not yet proved Observations 1 and 2. Another problem is that we only analyzed the optimality of the Whittle index if the problem satisfies the PCL-indexability with respect to some set family $\tilde{\Omega}$, but we have not yet explored other possible sets that may yield PCL to work as well. Furthermore, the strict calculation of Whittle indices for the entire belief state space is still open (even though we only need their order to be correct). All in all, more rooms for exploration of problems with infinite state spaces still await to be unlocked.

### v3

@@ 4 Conclusion
In this paper, we generalized the PCL framework for RMAB with finite state spaces to countable state spaces and propose an efficient algorithm to test the PCL-indexability and compute the Whittle index function for the partially observable RMAB with general observation models. As the number of belief states increases rapidly with time, the classic dynamic programming analysis for the RMAB problem suffers from the curse of dimensionality. Based on the achievable region and conservation law analysis studied in the literature for various RMAB problems, this paper is the first to extend the previous theoretical results in this direction to infinite-state scenarios by taking advantages of the theory of infinite-dimensional linear programming. Under a loose regularity assumption, we prove that the PCL-indexability is a sufficient condition for the Whittle indexability. Last, by introducing the concept of approximate state space, we build an algorithm to calculate the approximate Whittle index of each state using a finite-state adaptive greedy algorithm. Numerical results verified the strong performance of the Whittle index policy in our model.
Although we have quite a few theoretical analyses and have confirmed the effectiveness of the Whittle index algorithm, some problems remain to be solved in the future. An important one is that we only analyzed the optimality of the Whittle index under the condition that the problem satisfies PCL-indexability with respect to a specific set family $\tilde{\Omega}$, but we have not yet explored other possible sets that might also allow PCL to function effectively. Furthermore, the strict calculation of Whittle indices for the entire belief state space remains an open problem (even though only their relative order is required for optimal policy implementation). In summary, there are still many opportunities for further exploration of problems with infinite state spaces.